OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11 A. −2 B. −3 C. x can be any number as there are infinitely many solutions to the system. D. There is no x value as there is no solution to the system

OpenStudy (anonymous):

use elmination multiple top equation by 5 and the bottom equation by -6

OpenStudy (anonymous):

That will get u x....

OpenStudy (anonymous):

What's the answer?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

U multiply the first equation by 7 and the second by 3 and subtract them.

OpenStudy (anonymous):

the answer is -2 ?

OpenStudy (anonymous):

A IS THE ANSWER

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

Ok, i give up, it's -2

OpenStudy (anonymous):

U get -27y = 54 if u do what I said.

OpenStudy (amistre64):

3x − 6y = 3 7x − 5y = −11 y = (-33-21)/(-15+42) = - 54 / 27

OpenStudy (amistre64):

y = -2 :)